OpenStudy (anonymous):

What is the solution of the system? Use a graph. y = x + 5 y = –5x – 1

OpenStudy (anonymous):

ive already tried to do this got it wrong and i actually thought i had a good answer turns out i was wrong

OpenStudy (anonymous):

Graph them, the point where they intersect is your solution

OpenStudy (anonymous):

@CalebBeavers i did graph them i got the wrong answer

OpenStudy (anonymous):

@JakeV8 can you help me figure out how to do this ive already got it typed it in on my calculator

OpenStudy (tkhunny):

Graph them better! Let's follow the lines from the y-axis. I'll increment +1 in the y direction. y = x + 5 (0,5) ==> (1,6) ==> (2,7) y = -5x - 1 (0,-1) ==> (1,-6) ==> (2,-11) Hey, they're going the wrong direction! They're not going to intersect in this direction. Maybe we should look on the other side fo the y-axis! Let's decrement x by 1 at a time. y = x + 5 (0,5) ==> (-1,4) ==> (2,3) y = -5x - 1 (0,-1) ==> (-1,4) ==> (-2,9) Aha! What happened in the middle, there?!

OpenStudy (anonymous):

ummm ugh i have no idea what to do, i should probly cry myself to sleep for being so stupid

OpenStudy (tkhunny):

Now, none of that. Follow the points I enumerated. You will see the collision there at x = -1.

OpenStudy (anonymous):

i have it in my calculator but thats all i know how to do

OpenStudy (anonymous):

collision wait what? im in finding the collision or something else

OpenStudy (tkhunny):

Can you see the two lines?

OpenStudy (anonymous):

am i finding*

OpenStudy (anonymous):

yes i see where they come to together but im still lost kinda wanna give up

OpenStudy (tkhunny):

If you can get both lines on the screen, you should be able to pan and zoom to see the solution. You may need to spend a little more time in your owner's manual.

OpenStudy (anonymous):

:( i just messed up my whole graph on my calculator i think im done with this question im never going to understand this im sorry but thank you for trying to help

OpenStudy (tkhunny):

Hmmm...Really, I always find a new calculator confusing. Get you owner's manual out and work ALL the examples. Spend some quality time. What's your plan for the weekend? your calculator can be your friend, but not if you are not familiar with it. It takes a little time. Just take the time and don't be frustrated.

OpenStudy (anonymous):

@JakeV8 i dont even understand your graph :(

OpenStudy (anonymous):

im retaking this course because i already failed it 2 times freshman year and sophmore year :( i guess ill fail it for a third year now im a senior

OpenStudy (anonymous):

wow i just got reported for the second time for asking for the answer im not even asking for an answer i just want a similiar explaination

OpenStudy (anonymous):

@jennyjewell25 we can work this out... glad you aren't asking for answers :) since as we know, that's not really what the site is about... so don't worry about reports... you can still get this problem worked out.

OpenStudy (anonymous):

Here's your problem... it's a system of 2 equations, each with x and y... Equation 1: y = x + 5 Equation 2: y = –5x – 1 So, for graphing, let's look first at Equation # 1: it says, for any "x" value, the "y" value is the "x" value plus 5... do you see that? so, for x=0, y will be = 5... see?

OpenStudy (anonymous):

i understand it all but when you said it says, is where i got lost

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

"it says" just means that is what the first equation means... if you have an equation y = x + 5, that means, for any x value, to find the y value that goes with it, just add 5. So, if I am 5 years older than you, and my age is "y", and your age is "x", then... my age = your age + 5 or... y = x + 5 put your age in as "x", then add 5, and you get my age.

OpenStudy (anonymous):

ok i get that

OpenStudy (anonymous):

so that's all... the first equation is a line that shows y values that are 5 more than the x value at any particular point... if you ask "what is the y value at x = 2?" then the answer is, "y = x + 5" which is y = 2 + 5, or y = 7... so a point on that equation is (2,7)... the y value in the point is 5 more than the x value.

OpenStudy (anonymous):

i quit im sorry at this point id rather fail cause i deserve to fail for not understanding this when its explained to me cereval times